OpenStudy (anonymous):

Solve for x (x/sin0.6) + x = 20 The answer is 7.218 but I'm not so sure how

OpenStudy (lncognlto):

x/sin 0.6 + x = 20 Therefore x + sin 0.6x = 20(sin 0.6) therefore x(1 + sin 0.6) = 20(sin 0.6) tf x = 20(sin 0.6)/(1 + sin 0.6) tf x = 7.218 rounded...

OpenStudy (lncognlto):

I think I know this exact question, actually. I think answered it on a past exam paper the other day. xD

OpenStudy (anonymous):

I put that into my calculator and got nowhere close to 7.218...

OpenStudy (lncognlto):

It is 0.6 radians, not 0.6 degrees, so your calculator might be in the wrong mode. It comes out fine my side...

OpenStudy (anonymous):

Oh yeah thanks lol... Do you do AS math?

OpenStudy (lncognlto):

Yeah, I do, lol. I'm busy writing my final maths exams now. xD

OpenStudy (anonymous):

lol same cramming for the last few days

OpenStudy (lncognlto):

lol, me too. My next exam is next week, and then I'm finished with maths for school. ^-^ I'm in South Africa, by the way.

OpenStudy (anonymous):

I'm in New Zealand so we have our exam first... Could you help me with 7b of this paper please? http://papers.xtremepapers.com/CIE/Cambridge%20International%20A%20and%20AS%20Level/Mathematics%20(9709)/9709_s12_qp_11.pdf I'm not really sure what convergent means...

OpenStudy (lncognlto):

Hmm. Let me just get my maths book xD

OpenStudy (anonymous):

Ok thanks :D

OpenStudy (lncognlto):

Ok, well: when a series is convergent, the terms of the series decrease in magnitude and the sum to infinity of the series can be calculated by use of the formula a/(1-r), but only when l r l (the absolute value of r) is < 1. Therefore 0 < r < 1.

OpenStudy (lncognlto):

Now in the case of 7(b)(i),\[r = \frac{ 1 }{ 3 }\tan ^{2}\theta \]

OpenStudy (lncognlto):

therefore, \[0 \lt \frac{ 1 }{ 3 }\tan ^{2}\Theta \lt 1\]

OpenStudy (anonymous):

How exactly would I rearrange that?

OpenStudy (lncognlto):

Thus 0 < tan^2(theta) < 3 thus 0 < tan (theta) < sqr(3) therefore 0 < theta < tan^-1 sqr(3) therefore 0 < theta < pi/3

OpenStudy (lncognlto):

Or, in an equation, \[0 \lt \theta \lt \frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

I understand and for part two I did \[\frac{ 1 }{ 1-\frac{ 1 }{ 3 }\tan^{2}\times \frac{ 1 }{ 6 }pi }\] But my calculator keeps giving me the wrong value...

OpenStudy (lncognlto):

Hmm. Okay, let's try it by hand: \[\frac{ 1 }{ 1 - \frac{ 1 }{ 3 } \tan ^{2}\frac{ \pi }{ 6 }}\] Therefore, as \[\tan \frac{ \pi }{ 6 } = \frac{ 1 }{ \sqrt{3} }\] we get \[\frac{ 1 }{ 1 - \frac{ 1 }{ 3 }(\frac{ 1 }{ \sqrt{3} })^{2} }\]therefore \[\frac{ 1 }{ 1 - (\frac{ 1 }{ 3 })(\frac{ 1 }{ 3 }) }\] therefore\[\frac{ 1 }{ 1 - \frac{ 1 }{ 9 } }\] therefore \[\frac{ 1 }{ \frac{ 8 }{ 9 } }\] therefore \[S \infty = \frac{ 9 }{ 8 }\] or 1.125

OpenStudy (anonymous):

Wow thank you so so so so much for your help!!! I really appreciate you taking all this time and effort into helping me :D

OpenStudy (lncognlto):

No problem :) I finally understand that question myself now. I did it before, but it didn't make any sense to me then, lol.

OpenStudy (anonymous):

lol, well good luck for your exam!!!

OpenStudy (lncognlto):

Thanks, you too!

OpenStudy (anonymous):

Well see you later I need to go now, it was nice talking to you

OpenStudy (lncognlto):

Bye then :)